OpenStudy (tylermckinney16):

A salesperson sells $5780.00 worth of merchandise in one day. From working on commission his gross pay for that day was $179.18. What is his rate of commission?

OpenStudy (wolf1728):

5,780 * x = 179.18 x = 179.18/5,780 I think you can solve it fro there

OpenStudy (tylermckinney16):

$32.2580645161????is this right

OpenStudy (wolf1728):

You are calculating a commission rate and so your answer should be a percentage. x = 179.18/5,780 x = .031 .031*100 = 31%